# Exploring Mutual Empowerment Between Wireless Networks and RL-based LLMs: A Survey

### v2

@@ II-B Semantic Communication
In parallel with the wireless communication network infrastructure, many studies are also dedicated to developing the design of wireless communication, which can enhance the quality of transmission, make it faster, and address the current bottleneck of wireless communication. Traditional communication can provide good performance under low-noise environments, but failed communication is experienced when the channel noise exceeds a threshold. Semantic communication has been raised as a prominent pillar for the next wireless generation, where it considers the semantic meaning of the data in the transmission and eliminates the redundant information from the transmit signal. By jointly optimizing the source and channel encoders, the semantic features are extracted from the raw data and then compressed accordingly to the channel conditions, which improves the robustness against the physical noise and overcomes the “cliff effect”. The rise of deep learning (DL) has enabled semantic communication to differentiate the meaningful from the redundant and thus significantly reduce the number of symbols to be transmitted and save the bandwidth resource for other users [96, 97, 98, 99]. While the deep joint source-channel coding approach for semantic communication is mainly focused on data recovery, generative semantic communication offers freedom for wireless communication by allowing the reconstructed data at the receiver to be different from the original [100, 101, 102, 103]. However, it does not mean the two data are completely different; they actually share a semantic meaning. Specifically, in certain tasks such as image retrieval, instead of transmitting full-size image data, the generative AI semantic communication can convert them into descriptions, and the receiver will leverage the text description to obtain the corresponding images [104]. Additionally, the beauty of generative AI models is the ability to process multi-modality to improve the quality of the service with a low communication cost [105].
II-C Integrated Sensing and Communication
@@
Integrated Sensing and Communication (ISAC) is another interesting idea for tackling the wireless communication bottleneck by utilizing the same spectrum for both sensing and communication (avoiding latency time for channel switching) and the same hardware (antennas and transceivers). With radar sensing, we can actively implement beamforming to achieve the highest throughput and reduce communication latency. With the signal probe, ISAC provides high-accuracy environment sensing data, which leads to the improvement of channel state information (CSI) estimation. The accuracy of the CSI enables the improvement in modulation, coding, and finally, throughput rate and eases the wireless traffic [106]. In addition, ISAC not only can provide the current environment condition but also track the movement with high precision, which is ultimately useful for proactive beam steering and handover management [107, 108, 109]. Another advantage of user trajectory prediction is inter-user interference minimization by adjusting the resource allocation strategies[110, 111, 112]. The technique alone contains various benefits and can be fused together with other prominent directions, such as MEC, and semantic communication for further advancements [113, 114, 115, 116, 117]. Besides the ISAC technique, there exist numerous approaches to enhance the quality of the communication links, such as reconfigurable intelligent surface (RIS) to tune the signal reflection [118, 119, 120, 121, 122, 123], channel estimation [124, 125].
III Overview of RL-based LLMs
@@
Figure 4: Diagram depicting the three stages of integrating RL into training LLMs: (1) Supervised fine-tuning, (2) Reward definition, and (3) Policy optimization. Step 2 shows the comparison of three reward model definition approaches: RL from Human Feedback (RLHF), RL from AI Feedback (RLAIF), and Pure RL. RLHF and RLAIF share similar pipelines, differing only in how they generate rating responses.
The emergence of LLMs has unlocked immense potential for AI-driven innovations in wireless communication [126]. Within this domain, LLMs enhance network design, security, optimization, and management through their capabilities in data analysis, pattern recognition, and decision-making. These models find applications in diverse areas of wireless communication [127, 128, 129], including traffic allocation, failure detection, interference mitigation, and resource optimization, leading to substantial enhancements in network efficiency and reliability.
LLMs can be categorized into three common architectures: encoder-only, encoder-decoder, and decoder-only, stemming from the vanilla transformer model. While encoder-only models such as BERT [130] and DeBERTa [131] are proficient in analyzing network data logs and swiftly identifying irregularities, encoder-decoder architectures such as UL2 [132] and ChatGLM [36] are particularly effective for tasks requiring sequence-to-sequence transformations. These tasks include protocol conversion, optimization of signal processing, and enhancement of real-time communication capabilities [133, 2]. On the other hand, decoder-only architectures, such as DeepSeek-V3 [134], GPT-4o [135], are good at predictive modeling and automated troubleshooting. These models are highly effective at power allocation, spectrum sensing, and protocol understanding [126].
Their recent success is attributed to the efficient utilization of RL techniques in training the LLMs. RL techniques, including policy gradient methods [136], proximal policy optimization (PPO) [137] and group relative policy optimization (GRPO) [138], play a crucial role in optimizing LLMs’ decision-making processes in dynamic network environments. RL approaches in LLMs can be categorized into RL with Human Feedback (RLHF) [34], RL with Artificial Intelligence Feedback (RLAIF) [139], and Pure RL [28], each presenting distinct benefits in the training of LLMs for both general LLM tasks and more specialized wireless communication tasks. The overview of these approaches is illustrated in Fig. 4.
III-A Reinforcement Learning
RL stands out as a promising research area within machine learning that has gained substantial attention in recent years [33, 32, 39]. RL tasks involve learning how to solve a problem by taking a series of actions aimed at maximizing cumulative rewards [140]. To illustrate, we consider the classic game of Tic-Tac-Toe as shown in Fig. 5, where the objective is to form a row, column, or diagonal within a 3x3 grid before your opponent does. Success in this game hinges on making strategic moves (placing X’s or O’s in the grid), with each move influencing the outcome of subsequent actions. Imagine the player in this scenario as an AI agent; our goal is to train this AI agent to win the game. While traditional AI approaches, such as supervised and unsupervised learning, may struggle with this task due to its complexity, RL offers a fresh perspective. In RL, the problem is viewed through the lens of five key components: agent, environment, agent action, agent state, and rewards. Fig. 5 illustrates how these components interact to win the Tic-Tac-Toe game. In this approach, the agent acts based on the current situation, interacts with the environment to get feedback, and repeats this process until a winner is determined in each game, known as an episode in RL. The agent learns to improve its strategy by trial and error, aiming to maximize the rewards received in each episode.
@@ III-A Reinforcement Learning
Figure 5: An example of a reinforcement learning (RL) application in the Tic-Tac-Toe game. The agent learns optimal strategies through self-play, updating the state-action value function to enhance decision-making over time.
In NLP, we can view the agent as our LLM. The agent generates an action (response) based on its current state (prompt) and policy (LLM) to interact with the environment. The environment, in turn, responds to the agent’s action through rewards and transitions the current agent state to the next agent state (next prompt). Depending on our training objectives, the environment needs to be defined to provide rewards based on agent actions and states. This approach aims to maximize the number of rewards generated by the environment through our agent action. Specifically, in the context of LLM, defining the reward signal is crucial for optimizing the model using the RL approach. The difference in constructing reward models will be presented in the next section.
III-B RLHF-based LLMs
@@
RLHF [34] is an optimization method that integrates RL techniques with the reward model trained from the human feedback dataset to enhance the LLM’s ability to interpret and act according to the user’s intentions. By incorporating this human-generated reward model into the RL loop, RLHF strives to minimize the occurrence of inaccurate or uninformative outputs from the LLM. The common process of RLHF involves three common steps: supervised fine-tuning, reward model training, and policy optimization.
@@
Supervised Fine-Tuning: This step focuses on refining LLMs such as DeepSeek-V3 [134] and GPT-4o [135], transitioning them from a broad domain to one tailored for specific tasks. It involves collecting demonstration data and training a supervised policy to establish a robust foundation LLM for subsequent RL processes, as shown in step 1 of Fig. 4. Similar to the cold start phase of DeepSeek-R1 [28], this step aims to mitigate initial instability in RL training that may arise from the base model. The demonstration dataset includes pairs of prompt inputs and desired output labels provided by humans, which showcase the desired output behavior. For example, a prompt input might be ”Are there any specific locations where the signal drops significantly?” and the corresponding desired output would be the accurate locations where the signal is dropped. This data guides the model in learning the correct responses and behaviors for specific tasks.
@@
Reward Model Training: The reward model serves as the cornerstone of RL within LLM frameworks, providing a training signal for updating the model and guiding it toward desired behaviors. In the context of RLHF, this reward model assesses the LLM’s outputs by assigning them scalar values based on their quality. The primary objective of training in this framework is to optimize the reward obtained from the reward model for each interaction. To train the reward model, RLHF initially gathers outputs from multiple LLMs based on a given prompt. These outputs are then evaluated by a human labeler who ranks them according to specific criteria, ranging from best to worst. Subsequently, a supervised learning approach is employed to train the reward model, leveraging the labeled data to refine the model’s ability to evaluate and score the LLM outputs accurately, as shown in step 2 of Fig. 4. Several notable approaches, such as InstructGPT [34] and GPT-4 [26], adopt this strategy to guide their models effectively. By incorporating human feedback into the training process and iteratively refining the reward model’s evaluation capabilities, these advanced language models strive to produce outputs that better align with user expectations and preferences.
Policy Optimization: In RL, policy optimization is similar to model optimization in deep learning, involving updating model weights. However, a key distinction in RL is that the model does not receive immediate feedback at each step to indicate whether the update is moving in the correct or incorrect direction. Instead, the current policy generates outputs, which are subsequently evaluated by a reward model to generate corresponding reward scores. These reward scores are then utilized to update policy using policy optimization techniques such as PPO [137] and GRPO [138]. Consequently, the model receives feedback only after completing text generation or after a series of actions (steps), referred to as the ”delayed rewards” concept in RL.
III-C RLAIF-based LLMs
@@
RLAIF [139] follows a similar pipeline to RLHF, with the key distinction being that RLAIF offers an efficient approach to training a reward model by leveraging existing LLM. As illustrated in the Reward model training block of Step 2 of Fig. 4, RLAIF differs from RLHF by replacing human feedback labels with a process that involves presenting the same task to an independent, pre-trained LLM. RLAIF achieves this by selecting more harmless responses, such as multiple-choice options, generated within a predefined format that effectively conveys the target principles for the given task. Subsequently, log probability responses are computed to derive scalar scores, which serve as essential signals for the reward model’s training and optimization. In addition, varying the formats and principles employed in this process can lead to notably more robust model behavior. By adopting this strategy, the need for extensive human effort in dataset collection is reduced, as constructing datasets manually can be a time-intensive process.
By utilizing RLHF or RLAIF, LLMs can be adjusted to exhibit the intended output behavior. However, these approaches pose challenges related to computational overhead and inference time, thereby impeding the training progress of RL [141]. Additionally, the process of gathering data labels for training the reward model is laborious and expensive for both RLHF and RLAIF, rendering these approaches impractical for adaptation to different domains.
III-D Pure RL-based LLMs
Another innovative approach involves replacing the reward model with predefined reward rules, as proposed by DeepSeek-R1 [28], which has demonstrated exceptional performance in reasoning tasks for LLMs. Instead of relying on a pre-trained reward model derived from a rating dataset, DeepSeek-R1 implements a rule-based reward system comprising two primary types of rewards: accuracy rewards and format rewards. Accuracy rewards are designed to enable the model to assess the correctness of its responses. For instance, in scenarios such as math problems with deterministic solutions, the model can only earn a point if it provides the correct answer. On the other hand, format rewards force the model to respond in accordance with a predefined format to be considered a valid response and earn a point.

### v3

@@ 2.1.2. Towards Intelligent Network Architecture Design: The Role of Semantic Communication and ISAC
Semantic Communication. Traditional communication systems perform well in low-noise environments, but communication failures occur when channel noise exceeds a certain threshold (Trewavas, 2011). Semantic communication has emerged as a key technology for the next generation of wireless systems, as it focuses on the semantic meaning of the data being transmitted, eliminating redundant information from the transmitted signal to enhance efficiency (Raha et al., 2023c, b; Luo et al., 2022). By jointly optimizing the source and channel encoders, semantic features are extracted from the raw data and compressed based on the channel conditions. This approach enhances robustness against physical noise and mitigates the "cliff effect" (Getu et al., 2025). The rise of deep learning has enabled semantic communication to distinguish meaningful information from redundant data, significantly reducing the number of symbols transmitted and saving bandwidth resources for other users (Bourtsoulatze et al., 2019; Gündüz et al., 2023; Hassan et al., 2025; Yang et al., 2023a). While the deep joint source-channel coding approach for semantic communication primarily focuses on data recovery, generative semantic communication offers greater flexibility by allowing the reconstructed data at the receiver to differ from the original (Grassucci et al., 2024; Xu et al., 2024a; Nguyen et al., 2025b; Fan et al., 2024). However, this does not imply that the two sets of data are entirely different; they still share the same semantic meaning. Specifically, in tasks such as image retrieval, instead of transmitting full-sized image data, generative AI-based semantic communication can convert the images into descriptions. The receiver then uses these text descriptions to retrieve the corresponding images (Yang et al., 2024b). Additionally, the advantage of generative AI models lies in their ability to process multimodal data, enhancing service quality while maintaining low communication costs (Zhao et al., 2024b).
@@
Integrated Sensing and Communication. ISAC (Zhu et al., 2024b; Cang et al., 2024) offers an innovative solution to address the wireless communication bottleneck by utilizing the same spectrum for both sensing and communication, thereby eliminating the latency associated with channel switching, as well as the same hardware (antennas and transceivers). Through radar sensing, ISAC enables active beamforming to achieve optimal throughput and reduce communication latency. Additionally, with signal probing, ISAC provides high-accuracy environmental sensing data, which enhances channel state information (CSI) estimation. The improved CSI accuracy leads to better modulation and coding, ultimately increasing throughput and alleviating wireless traffic (Kim et al., 2023). In addition, ISAC not only provides real-time environmental conditions but also tracks movement with high precision, making it highly beneficial for proactive beam steering and handover management (Liu et al., 2022; Wei et al., 2022a; Xie et al., 2024). Another advantage of user trajectory prediction is the minimization of inter-user interference through optimized resource allocation strategies (Niu et al., 2024b; Wang et al., 2024a; Niu et al., 2024a). This technique offers various benefits on its own and can be integrated with other advanced technologies, such as mobile edge computing (MEC) and semantic communication, to drive further advancements (Cang et al., 2024; Liu et al., 2023; Lu et al., 2024a; Zhu et al., 2024b; Yang et al., 2024a). In addition to ISAC, numerous other approaches, such as channel estimation, also exist to enhance the quality of communication links (Kim and Hong, 2022; Kim et al., 2024; Soltani et al., 2019).
Wireless Security. The explosive increase in the number of connected devices induces security threats to the users, as eavesdroppers and jammers have a higher chance of interfering with the communication link due to the density of the wireless network (Zhao et al., 2025a). To address this, various types of wireless security have been studied, such as the physical layer, network layer and application layer, in which the physical layer security (PLS) has drawn the attention of the research community due to its massive potential. Specifically, PLS takes advantage of the unique physical characteristics of hardware devices and channels for authentication (Afeef et al., 2022). Additionally, the development of AI technologies has been integrated into PLS, which successfully resolves the complexity problem of the environment and channels. In the work (Zhao et al., 2024a), the authors have demonstrated the benefits and drawbacks of traditional AI; then, they illustrate the merit of the adoption of generative AI for a secure physical layer.
2.1.3. Towards Seamless Connectivity: The Role of NTNs and MEC
To meet the growing demands of high-quality service and seamless connectivity, modern wireless networks are undergoing a shift towards the next-generation computing and communication paradigm (Hu et al., 2015; Liu et al., 2021). This shift aims to enable more intelligent, efficient, and context-aware services across diverse environments. Key technological enablers of this trend include non-terrestrial networks (NTNs) (Nguyen et al., 2023b), which offer ubiquitous connectivity beyond terrestrial infrastructure, and mobile edge computing (MEC), which brings computation closer to the user to reduce latency and alleviate core network congestion. Additionally, reconfigurable intelligent surfaces (RIS) enhance signal propagation by dynamically shaping the wireless environment, while emerging high-frequency communication technologies such as millimeter-wave (mmWave) and terahertz (THz) enable ultra-high-speed data transmission. Massive multiple-input multiple-output (MIMO) (Cho et al., 2010), leveraging large antenna arrays, further contributes to this integration by supporting concurrent high-throughput links for numerous users. Collectively, these technologies form the foundation of a new network paradigm in which computing and communication are no longer isolated components, but deeply intertwined to support future intelligent services.
Non-Terrestrial Networks. The existing wireless communication infrastructure relies heavily on fixed base stations, which struggle to deliver high-quality services in rural areas and special scenarios such as natural disaster response and military operations (Imran et al., 2024). To meet the growing demand for ubiquitous connectivity, NTNs are designed to provide coverage across vast areas (Nguyen et al., 2024a, 2023b; Hassan et al., 2023). These networks can be deployed using unmanned aerial vehicles (UAVs) (Nguyen et al., 2023b), high-altitude platforms (HAPs) (Tozer and Grace, 2001), or satellites operating in low-Earth orbit (LEO) (Talgat et al., 2024), medium-Earth orbit (MEO) (Marshall et al., 2021), and geostationary orbit (GEO) (Talgat et al., 2024). The NTNs can be established by deploying a single type of these devices or a combination of multiple types, each offering distinct advantages. UAVs offer flexibility for the network based on ease of flight path adjustment, while satellites significantly expand coverage and offer greater resources in terms of computing power and energy compared to UAVs. With the moving property, the NTNs can create communication connections to remote users beyond the reach of fixed base stations. However, the proposal of NTNs, along with the MEC, increases the complexity of the wireless communication infrastructure, and it needs to be managed efficiently to be able to exploit the proposal’s capabilities fully (Chen et al., 2024b; Mao et al., 2017c, b).
Massive Machine Type Communications. The concept of smart factories and smart cities has flourished in recent years with the assistance of automated devices such as sensors, tracking devices, meters, and actuators. Communication between these devices is typically carried out among intelligent machines with minimal human intervention—or ideally, none at all—which is referred to as machine-type communications (MTC) (Dutkiewicz et al., 2017). On the other hand, “massive MTC” refers to scenarios where the number of connected devices is at least ten times greater than current levels, with varying quality-of-service (QoS) requirements imposed on the cellular network (Dawy et al., 2016; Bockelmann et al., 2016). The technique has paved the way for the smart environment and is expected to be the pillar of next-generation communication (Sharma and Wang, 2020). However, the key open questions about large-scale MTC deployment over cellular networks are when it will occur and how it should be designed and planned.
Beamforming for mmWave and THz Communication. With the large bandwidth offered by mmWave (30-300 GHz) (Wang et al., 2018) and THz (0.1-10 THz) (Akyildiz et al., 2014) communication, the wireless system can facilitate more users with ultra-high-speed data and low interference. These technologies can address the spectrum scarcity associated with the exponential growth of connected devices. However, since the wavelength of both techniques is significantly short, the short wavelength can be easily reflected, absorbed, or blocked by obstacles in the communication link. Therefore, it can only operate within short-range distances, and this is where the beamforming technique steps in to provide steering ability and create narrow beams with concentrated power toward the user, thus improving SNR and enabling a more extended communication range (Ning et al., 2023; Zhou et al., 2025). Additionally, in a multiuser scenario, beamforming can achieve secure communication by preventing both active and passive eavesdroppers in the wireless transmission process with an appropriate resource allocation algorithm (Ng et al., 2014).
@@ 2.1.3. Towards Seamless Connectivity: The Role of NTNs and MEC
Figure 3. Diagram depicting the three stages of integrating RL into training LLMs: (1) Supervised fine-tuning, (2) Reward definition, and (3) Policy optimization (Schulman et al., 2017; Shao et al., 2024d). Step 2 shows the comparison of three reward model definition approaches: RL from Human Feedback (RLHF) (Ouyang et al., 2022), RL from AI Feedback (RLAIF) (Bai et al., 2022), and Pure RL (Guo et al., 2025). RLHF and RLAIF share similar pipelines, differing only in how they generate rating responses.
Mobile Edge Intelligence. MEI (Hu et al., 2015) is an emerging technology that enhances network performance by providing computational capabilities closer to end-users. These devices are widely distributed and positioned closer to users than cloud servers, thereby reducing the communication bottleneck for cloud servers (Qiao et al., 2023b). By deploying computing resources closer to end-users, MEI minimizes communication delays and alleviates network congestion, making it particularly suitable for real-time applications (Qiao et al., 2024d). Additionally, by decoupling the data and control planes into individuals, the SDN (Haque and Abu-Ghazaleh, 2016) can control the network effectively by adjusting the traffic routing, spectrum allocation to multiple access points, and slicing networks for different applications. Similar to MEI, another emerging technology in wireless communication is RIS (Liu et al., 2021; Huang et al., 2019). RIS is designed to establish an indirect communication link between the transmitter and receiver by dynamically controlling signal reflection. This technology is particularly beneficial in environments where direct line-of-sight communication is obstructed (Aung et al., 2024b, a). RIS has been deployed in various scenarios, including smart buildings, vehicular networks, UAVs, and high-rise structures (Park et al., 2024; Aung et al., 2023; Hassan et al., 2024b). By intelligently adjusting reflection coefficients, RIS can optimize signal propagation, reduce interference, and improve wireless link reliability. The integration of RIS with MEI and SDN further enhances network adaptability, providing a promising solution for future wireless communication systems. In addition to the substantial efforts dedicated to optimizing wireless networks, the rapid advancement of large language models (LLMs) has demonstrated significant potential in enabling AI-driven solutions for intelligent network management and optimization (Xu et al., 2024b, 2025b). The following subsection first introduces the core technologies underlying LLMs, with representative examples such as DeepSeek, and then explores the emerging synergy between wireless networks and LLMs.
2.2. Key Techniques Behind DeepSeek and Beyond
The emergence of LLMs has unlocked immense potential for AI-driven innovations in wireless communication (Shao et al., 2024c). These models play a transformative role in network design, security, optimization, and management by leveraging their advanced capabilities in data analysis, pattern recognition, and decision-making. LLMs are applied across various aspects of wireless communication (Zhang et al., 2024a; Qiu et al., 2024a; Jiang et al., 2024), including traffic allocation, failure detection, interference mitigation, and resource optimization, leading to substantial enhancements in network efficiency and reliability. LLMs can be categorized into three common architectures: encoder-only, encoder-decoder, and decoder-only, stemming from the vanilla transformer model. While encoder-only models such as BERT (Devlin et al., 2019) and DeBERTa (He et al., 2021) are proficient in analyzing network data logs and swiftly identifying irregularities, encoder-decoder architectures such as UL2 (Tay et al., 2022) and ChatGLM (GLM et al., 2024) are particularly effective for tasks requiring sequence-to-sequence transformations. These tasks include protocol conversion, optimization of signal processing, and enhancement of real-time communication capabilities (Verma and Pilanci, 2024; Nguyen et al., 2024c). On the other hand, decoder-only architectures, such as DeepSeek-V3 (Liu et al., 2024b), GPT-4o (Hurst et al., 2024), are good at predictive modeling and automated troubleshooting. These models are highly effective at power allocation, spectrum sensing, and protocol understanding (Shao et al., 2024c). Their recent success is attributed to the efficient utilization of reinforcement learning (RL) techniques in training the LLMs. RL techniques, including policy gradient methods (Wang and Zou, 2022), proximal policy optimization (PPO) (Schulman et al., 2017) and group relative policy optimization (GRPO) (Shao et al., 2024d), play a crucial role in optimizing LLMs’ decision-making processes in dynamic network environments. RL-based training strategies for LLMs can be broadly categorized into three main approaches: RL with human feedback (RLHF) (Ouyang et al., 2022), RL with artificial intelligence feedback (RLAIF) (Bai et al., 2022), and Pure RL (Guo et al., 2025). Each of these methods offers distinct advantages in the training of LLMs for both general LLM tasks and more specialized wireless communication tasks (Xu et al., 2024b, 2025b). An overview of these approaches is illustrated in Fig. 3.
@@ 2.2. Key Techniques Behind DeepSeek and Beyond
Figure 4. An example of an RL application in the Tic-Tac-Toe game. The agent learns optimal strategies through self-play, updating the state-action value function to enhance decision-making over time.
2.2.1. Reinforcement Learning
RL stands out as a promising research area within machine learning that has gained substantial attention in recent years (Wang et al., 2024e; Milani et al., 2024; Lei et al., 2020). RL tasks involve learning how to solve a problem by taking a series of actions aimed at maximizing cumulative rewards (Sutton and Barto, 2018). To illustrate, we consider the classic game of Tic-Tac-Toe as shown in Fig. 4, where the objective is to form a row, column, or diagonal within a 3x3 grid before your opponent does. Success in this game hinges on making strategic moves (placing X’s or O’s in the grid), with each move influencing the outcome of subsequent actions. Imagine the player in this scenario as an AI agent; our goal is to train this AI agent to win the game. While traditional AI approaches, such as supervised and unsupervised learning, may struggle with this task due to its complexity, RL offers a fresh perspective. In RL, the problem is viewed through the lens of five key components: agent, environment, agent action, agent state, and rewards. Fig. 4 illustrates how these components interact to win the Tic-Tac-Toe game. In this approach, the agent acts based on the current situation, interacts with the environment to get feedback, and repeats this process until a winner is determined in each game, known as an episode in RL. The agent learns to improve its strategy by trial and error, aiming to maximize the rewards received in each episode. In natural language processing (NLP), we can view the agent as our LLM. The agent generates an action (response) based on its current state (prompt) and policy (LLM) to interact with the environment. The environment, in turn, responds to the agent’s action through rewards and transitions the current agent state to the next agent state (next prompt). Depending on our training objectives, the environment needs to be defined to provide rewards based on agent actions and states. This approach aims to maximize the number of rewards generated by the environment through our agent action. Specifically, in the context of LLM, defining the reward signal is crucial for optimizing the model using the RL approach. The difference in constructing reward models will be presented in the next section.
2.2.2. RLHF-based LLMs
@@
RLHF (Ouyang et al., 2022) is an optimization method that integrates RL techniques with the reward model trained from the human feedback dataset to enhance the LLM’s ability to interpret and act according to the user’s intentions. By incorporating this human-generated reward model into the RL loop, RLHF strives to minimize the occurrence of inaccurate or uninformative outputs from the LLM. The common process of RLHF involves three common steps: supervised fine-tuning, reward model training, and policy optimization.
@@
Supervised Fine-Tuning. This step focuses on refining LLMs such as DeepSeek-V3 (Liu et al., 2024b) and GPT-4o (Hurst et al., 2024), transitioning them from a broad domain to one tailored for specific tasks. It involves collecting demonstration data and training a supervised policy to establish a robust foundation LLM for subsequent RL processes, as shown in step 1 of Fig. 3. Similar to the cold start phase of DeepSeek-R1 (Guo et al., 2025), this step aims to mitigate initial instability in RL training that may arise from the base model. The demonstration dataset includes pairs of prompt inputs and desired output labels provided by humans, which showcase the desired output behavior. For example, a prompt input might be "Are there any specific locations where the signal drops significantly?" and the corresponding desired output would be the accurate locations where the signal is dropped. This data guides the model in learning the correct responses and behaviors for specific tasks.
Figure 5. An example of CoT. Different from standard prompting, CoT prompting explicitly outlines the reasoning process, leading to more interpretable and accurate results.
@@
Reward Model Training. The reward model serves as the cornerstone of RL within LLM frameworks, providing a training signal for updating the model and guiding it toward desired behaviors. In the context of RLHF, this reward model assesses the LLM’s outputs by assigning them scalar values based on their quality. The primary objective of training in this framework is to optimize the reward obtained from the reward model for each interaction. To train the reward model, RLHF initially gathers outputs from multiple LLMs based on a given prompt. These outputs are then evaluated by a human labeler who ranks them according to specific criteria, ranging from best to worst. Subsequently, a supervised learning approach is employed to train the reward model, leveraging the labeled data to refine the model’s ability to evaluate and score the LLM outputs accurately, as shown in step 2 of Fig. 3. Several notable approaches, such as InstructGPT (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023), adopt this strategy to guide their models effectively. By incorporating human feedback into the training process and iteratively refining the reward model’s evaluation capabilities, these advanced language models strive to produce outputs that better align with user expectations and preferences.
Policy Optimization. In RL, policy optimization shares similarities with model optimization in deep learning, as both involve updating model weights. However, RL differs in that the model does not receive immediate feedback at each step to indicate whether the update direction is correct. Instead, the current policy produces outputs that are later assessed by a reward model, generating corresponding reward scores. This delayed feedback mechanism, known as "delayed rewards" in RL, means the model only receives evaluations after completing text generation or a sequence of actions (steps). Optimization in RL is commonly performed using methods like PPO (Schulman et al., 2017) or GRPO (Shao et al., 2024d). PPO, a widely adopted approach in RL, ensures stability during policy updates but requires both policy and value models, which can lead to high computational costs for large-scale tasks. In contrast, GRPO introduces a novel group-relative optimization framework that emphasizes relative performance within learning groups rather than absolute metrics, reducing computational overhead while maintaining effectiveness.
@@
RLAIF (Bai et al., 2022) follows a similar pipeline to RLHF, with the key distinction being that RLAIF offers an efficient approach to training a reward model by leveraging existing LLM. As illustrated in the reward model training block of step 2 of Fig. 3, RLAIF differs from RLHF by replacing human feedback labels with a process that involves presenting the same task to an independent, pre-trained LLM. RLAIF achieves this by selecting more harmless responses, such as multiple-choice options, generated within a predefined format that effectively conveys the target principles for the given task. Subsequently, log probability responses are computed to derive scalar scores, which serve as essential signals for the reward model’s training and optimization. In addition, varying the formats and principles employed in this process can lead to notably more robust model behavior. By adopting this strategy, the need for extensive human effort in dataset collection is reduced, as constructing datasets manually can be a time-intensive process. By utilizing RLHF or RLAIF, LLMs can be adjusted to exhibit the intended output behavior. However, these approaches pose challenges related to computational overhead and inference time, thereby impeding the training progress of RL (Cao et al., 2024). Additionally, the process of gathering data labels for training the reward model is laborious and expensive for both RLHF and RLAIF, rendering these approaches impractical for adaptation to different domains.
2.2.3. Pure RL-based LLMs